OpenStudy (anonymous):

Pleaasee helppp Let f and g be two functions defined by f(x)=x+5 g(x)=(x-1)(x+3) for any real number x . Find f/g(-4) and then find all values that are NOT in the domain of f/g(-4)

OpenStudy (anonymous):

f(x)=x+5 g(x)=(x-1)(x+3) ; g(x)=x^2+3x-x-3; g(x)=x^2+2x-3

OpenStudy (anonymous):

You know how to foil, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(4)=(4)+5 g(4)=(4)^2+2(4)-3 f(4)=9 g(4)=21. 9/21=(3/7)

OpenStudy (anonymous):

(x+5)/(x^2+2x-3) The domain is all real numbers except x=-3, and x=1

OpenStudy (anonymous):

Using x=-3 or x=1 will give you undefined

OpenStudy (anonymous):

oh okkaay thank you!